# Stronger Constraints on Primordial Black Holes as Dark Matter Derived from the Thermal Evolution of the Intergalactic Medium over the Last Twelve Billion Years

### v1

@@ VIII Conclusions
We have self-consistently modeled the temperature and ionization histories of the IGM, incorporating both astrophysical reionization and PBH energy injection. By comparing our predictions with Lyman-$\alpha$ forest temperature measurements, we have placed the second strongest constraints on PBH abundance in the mass range $7\times 10^{15}$ g to $2\times 10^{17}$ g, improving upon previous IGM temperature-based limits by approximately an order of magnitude. This improvement arises from our inclusion of lower-redshift data, where PBH-induced heating via Coulomb interactions is most significant, as well as our extension of the analysis to redshift zero, full incorporation of He iii evolution, and use of a physically motivated reionization model. We have also slightly expanded the exploration of PBH parameter space to the lower end of the mass range ($10^{17}\,\mathrm{g}$ to $10^{22}\,\mathrm{g}$), where no prior constraints were available (see, however, recent proposals [49, 85, 86, 87, 88, 89] to cover this mass window). While the strongest constraints on PBHs currently come from the 511 keV gamma-ray line, they rely on complex and uncertain astrophysical modeling, including positron propagation in Galactic magnetic fields, and dark matter density profiles. In contrast, our IGM temperature constraints are based on well-understood cosmological heating and cooling processes, independent of local PBH clustering and astrophysical uncertainties. Since PBH energy injection has the strongest impact at low redshifts, more precise future measurements of the IGM temperature at low redshifts (e.g., mock $T_{\mathrm{IGM}}$ data at $z=0.1$ from [90]) could further improve these constraints. More broadly, this study establishes a framework for leveraging the IGM thermal history to constrain a wide range of exotic energy injection scenarios, providing a powerful probe for testing beyond-Standard Model physics, including various dark matter models.
IX Acknowledgments
@@ A.1 Temperature Evolution
×$\quad\times\sum_{X}\left(\mathcal{H}^{\mathrm{ion}}_{eX}+\mathcal{H}_{X}^{\mathrm{rec}}+\mathcal{H}_{eX}^{\mathrm{exc}}+\mathcal{H}^{\mathrm{br}}\right).$ (20)
The summation runs over relevant species: $X=\mathrm{H}\,\textsc{i},\mathrm{He}\,\textsc{i},\mathrm{He}\,\textsc{ii}$ for collisional ionization and excitation, and $X=\mathrm{H}\,\textsc{ii},\mathrm{He}\,\textsc{ii},\mathrm{He}\,\textsc{iii}$ for recombination and Bremsstrahlung cooling. The cooling rates $\mathcal{H}$ are computed following Appendix B4 of Ref. [78]. Photoheating from reionization is given by:
@@
×$\quad\times n_{\mathrm{H}}(x_{\mathrm{HI}}\varepsilon_{\mathrm{HI}}+x_{\mathrm{HeI}}\varepsilon_{\mathrm{HeI}}+x_{\mathrm{HeII}}\varepsilon_{\mathrm{HeII}}).$ (21)
The photoheating rates $\varepsilon$ are taken from Table D1 of Ref. [77].
@@ A.2 Ionization Evolution
where $\Gamma_{e\mathrm{HI}}^{\mathrm{ion}}$ is the collisional ionization rate and $\alpha_{A,\mathrm{HI}}$ is the Case-A recombination coefficient, both taken from Appendix B4 of Ref. [78]. The contribution from reionization is:
@@
where $\Gamma_{\gamma\mathrm{HI}}$ is the photoionization rate of H i, taken from Table D1 of Ref. [77]. For He ii ionization:
@@
where He ii increases due to the ionization of He i and recombination of He iii, and decreases due to the ionization of He ii and recombination of He ii. The collisional ionization rates and Case-A recombination coefficients are taken from Appendix B4 of Ref. [78]. The reionization contribution to He ii is:
@@
where the photoionization rates are taken from Table D1 of Ref. [77].
@@ Appendix B Coulomb Heating and Energy Deposition
High-energy electrons and positrons injected by PBHs deposit energy into the IGM through various processes. After reionization, the dominant energy deposition channel is Coulomb heating, where energetic electrons transfer their kinetic energy to background free electrons, increasing the IGM temperature. Additionally, injected energy is also distributed into collisional ionization, collisional excitation, and inverse Compton scattering. In this work, we follow the methodology of Ref. [76] to model the heating due to PBH-injected electrons and positrons. This section specifically focuses on how we recomputed $f_{\rm heat}(z,{\textbf{x}})$ (as defined in Equation 9 in the main text) that encodes the fraction of injected energy which converts to heat.
@@ B.1 Coulomb Heating in the IGM
[…]
Here, we consider $N$ energy bins labeled $10,1,2,3,...,N-$ 1. All elements below the diagonal are zero, ensuring that electrons can only lose energy.
@@ B.2 Recursive Energy Deposition Framework
Let $R_{\mathrm{heat}}(E_{n})$ represent the total energy deposited into Coulomb heating, accounting for the cumulative energy transferred from the injected electron until its energy is fully dissipated. Over the short time interval $\Delta t$, the injected electron undergoes various cooling processes, including Coulomb interactions, collisional ionization, collisional excitation, and inverse Compton scattering. These interactions produce a secondary electron spectrum $\mathrm{d}\textbf{N}_{\mathrm{sec}}/\mathrm{d}\textbf{E}$ (see Equations 32 and 33 for the case considering only Coulomb heating, and Ref. [75] for cross sections and secondary spectra associated with other processes). A fraction of the injected electron’s energy, $\Delta E_{\mathrm{heat}}$ (given by Equation 26), is promptly deposited as heat within the same time interval $\Delta t$. The remaining energy is carried by the secondary electron, which subsequently loses energy through Coulomb interactions, described by $R_{\mathrm{heat}}(E)$, where $E$ represents the energy of the secondary electron. This process leads to the following recursive relation:
@@
Appendix C Table of Constraints on $f_{\mathrm{PBH}}$
We present the numerical values of the upper limits on $f_{\mathrm{PBH}}$ derived from our analysis. The table below lists the constraints for different PBH masses, ranging from $1\times 10^{15}$ g to $2\times 10^{17}$ g.

### v2

@@ VIII Conclusions
We have self-consistently modeled the temperature and ionization histories of the IGM, incorporating both astrophysical reionization and PBH energy injection. By comparing our predictions with Lyman-$\alpha$ forest temperature measurements, we have placed the second strongest constraints on PBH abundance in the mass range $7\times 10^{15}$ g to $2\times 10^{17}$ g, improving upon previous IGM temperature-based limits by approximately an order of magnitude. This improvement arises from our inclusion of lower-redshift data, where PBH-induced heating via Coulomb interactions is most significant, as well as our extension of the analysis to redshift zero, full incorporation of He iii evolution, and use of a physically motivated reionization model. We have also slightly expanded the exploration of PBH parameter space to the lower end of the mass range ($10^{17}\,\mathrm{g}$ to $10^{22}\,\mathrm{g}$), where no prior constraints were available (see, however, recent proposals [Ray:2021mxu, Esser:2022owk, Tamta:2024pow, Jung:2019fcs, Gawade:2023gmt, Fedderke:2024wpy] to cover this mass window). While the strongest constraints on PBHs currently come from the 511 keV gamma-ray line, they rely on complex and uncertain astrophysical modeling, including positron propagation in Galactic magnetic fields, and dark matter density profiles. In contrast, our IGM temperature constraints are based on well-understood cosmological heating and cooling processes, independent of local PBH clustering and astrophysical uncertainties. Since PBH energy injection has the strongest impact at low redshifts, more precise future measurements of the IGM temperature at low redshifts (e.g., mock $T_{\mathrm{IGM}}$ data at $z=0.1$ from [2022MNRAS.515.2188H]) could further improve these constraints. More broadly, this study establishes a framework for leveraging the IGM thermal history to constrain a wide range of exotic energy injection scenarios, providing a powerful probe for testing beyond-Standard Model physics, including various dark matter models.
Acknowledgments
@@ A.1 Temperature Evolution
$\quad\times\sum_{X}\left(\mathcal{H}^{\mathrm{ion}}_{eX}+\mathcal{H}_{X}^{\mathrm{rec}}+\mathcal{H}_{eX}^{\mathrm{exc}}+\mathcal{H}^{\mathrm{br}}\right).$ (20)
The summation runs over relevant species: $X=\mathrm{H}\,\textsc{i},\mathrm{He}\,\textsc{i},\mathrm{He}\,\textsc{ii}$ for collisional ionization and excitation, and $X=\mathrm{H}\,\textsc{ii},\mathrm{He}\,\textsc{ii},\mathrm{He}\,\textsc{iii}$ for recombination and Bremsstrahlung cooling. The cooling rates $\mathcal{H}$ are computed following Appendix B4 of Ref. [2007MNRAS.374..493B]. Photoheating from reionization is given by:
@@
$\quad\times n_{\mathrm{H}}(x_{\mathrm{HI}}\varepsilon_{\mathrm{HI}}+x_{\mathrm{HeI}}\varepsilon_{\mathrm{HeI}}+x_{\mathrm{HeII}}\varepsilon_{\mathrm{HeII}}).$ (21)
The photoheating rates $\varepsilon$ are taken from Table D1 of Ref. [2019MNRAS.485...47P].
@@ A.2 Ionization Evolution
where $\Gamma_{e\mathrm{HI}}^{\mathrm{ion}}$ is the collisional ionization rate and $\alpha_{A,\mathrm{HI}}$ is the Case-A recombination coefficient, both taken from Appendix B4 of Ref. [2007MNRAS.374..493B]. The contribution from reionization is:
@@
where $\Gamma_{\gamma\mathrm{HI}}$ is the photoionization rate of H i, taken from Table D1 of Ref. [2019MNRAS.485...47P]. For He ii ionization:
@@
where He ii increases due to the ionization of He i and recombination of He iii, and decreases due to the ionization of He ii and recombination of He ii. The collisional ionization rates and Case-A recombination coefficients are taken from Appendix B4 of Ref. [2007MNRAS.374..493B]. The reionization contribution to He ii is:
@@
where the photoionization rates are taken from Table D1 of Ref. [2019MNRAS.485...47P].
@@ Appendix B Coulomb Heating and Energy Deposition
High-energy electrons and positrons injected by PBHs deposit energy into the IGM through various processes. After reionization, the dominant energy deposition channel is Coulomb heating, where energetic electrons transfer their kinetic energy to background free electrons, increasing the IGM temperature. Additionally, injected energy is also distributed into collisional ionization, collisional excitation, and inverse Compton scattering. In this work, we follow the methodology of Ref. [2020PhRvD.101b3530L] to model the heating due to PBH-injected electrons and positrons. This section specifically focuses on how we recomputed $f_{\rm heat}(z,{\textbf{x}})$ (as defined in Equation 9 in the main text) that encodes the fraction of injected energy which converts to heat.
@@ B.1 Coulomb Heating in the IGM
$P_{n}$ $=1-\frac{\Delta E_{\mathrm{heat}}}{E_{n}-E_{n-1}},\quad\text{(diagonal elements)},$ (34)
$P_{n-1,n}$ $=\frac{\Delta E_{\mathrm{heat}}}{E_{n}-E_{n-1}},\quad\text{(upper-diagonal elements)}.$ (35)
Here, we consider $N$ energy bins labeled 0,1,2,3,,N10,1,2,3,...,N-1. All elements below the diagonal are zero, ensuring that electrons can only lose energy.
@@ B.2 Recursive Energy Deposition Framework
Let $R_{\mathrm{heat}}(E_{n})$ represent the total energy deposited into Coulomb heating, accounting for the cumulative energy transferred from the injected electron until its energy is fully dissipated. Over the short time interval $\Delta t$, the injected electron undergoes various cooling processes, including Coulomb interactions, collisional ionization, collisional excitation, and inverse Compton scattering. These interactions produce a secondary electron spectrum $\mathrm{d}\textbf{N}_{\mathrm{sec}}/\mathrm{d}\textbf{E}$ (see Equations 32 and 33 for the case considering only Coulomb heating, and Ref. [2009PhRvD..80d3526S] for cross sections and secondary spectra associated with other processes). A fraction of the injected electron’s energy, $\Delta E_{\mathrm{heat}}$ (given by Equation 26), is promptly deposited as heat within the same time interval $\Delta t$. The remaining energy is carried by the secondary electron, which subsequently loses energy through Coulomb interactions, described by $R_{\mathrm{heat}}(E)$, where $E$ represents the energy of the secondary electron. This process leads to the following recursive relation:
@@
Figure 5: Constraints on the PBH fraction $f_{\mathrm{PBH}}$ as a function of PBH mass $M_{\mathrm{PBH}}$. Our new constraint (red solid line) is shown alongside existing bounds from: extragalactic gamma rays [Carr:2009jm] in blue, the 511 keV gamma-ray line [2024PhRvD.110l3022L] in orange, modifications of the CMB spectrum [Clark:2016nst] in green, electron/positron measurements from Voyager 1 [Boudaud:2018hqb] in cyan, the MeV background [Laha:2020ivk] in dodger blue, heating effects in the dwarf galaxy Leo T [Laha:2020vhg] in brown, neutrino evaporation limits from Super-Kamiokande [Dasgupta:2019cae] in maroon, COMPTEL GC observations [Coogan:2020tuf] in red, and the X-ray background [2024PhRvD.110l3022L] in purple. Some of these constraints, including the 511 keV one (orange), depends on assumed dark matter density profile and can therefore be relaxed under different profile choices, whereas our constraint is based on the cosmological dark matter density and is independent of such uncertainties.
Appendix C Table of Constraints on $f_{\mathrm{PBH}}$
We present the numerical values of the upper limits on $f_{\mathrm{PBH}}$ derived from our analysis. The table below (see Table 1) lists the constraints for different PBH masses, ranging from $1\times 10^{15}$ g to $2\times 10^{17}$ g.
@@ Appendix C Table of Constraints on $f_{\mathrm{PBH}}$
Appendix D Comparison with Existing Constraints on $f_{\mathrm{PBH}}$
In this appendix, we present a comprehensive comparison of our newly derived constraint on $f_{\mathrm{PBH}}$ with various existing bounds reported in the literature. Figure 5 displays these constraints as a function of $M_{\mathrm{PBH}}$ for a monochromatic mass function of PBHs.